OpenStudy (anonymous):

The soccer team has lost 3 out of 7 games so far this season. Approximately what percent of games has the team won? a,30% b,40% c,43% d,57% please help me

OpenStudy (anonymous):

3/7= .48 = 48%

OpenStudy (anonymous):

its not on the options

OpenStudy (anonymous):

Well we know they lost 3/7 games, so they must have won the other four (as no ties are mentioned), therefore: 7-3=4. With this, we can say that, the number of games won over the total number gives us what we wish: \[ \frac{4}{7}\approx .57142 \]

OpenStudy (anonymous):

wait i dont get it

OpenStudy (anonymous):

So, we have a total of 7 games, and we know the soccer team has lost 3 of them. We know that the soccer team can either lose the game or win it, which means that the total number of games won added to the total number of games lost must be the total number of games. We have that the percentage is the ratio between the number of games that have been won, and the total number of games played, so, with this, we get that 7-3=4 games won, and the total number of games is 7, so, 4/7=.51742=57.1%

OpenStudy (anonymous):

thanks so much